OpenStudy (anonymous):

Factor these two questions fully please: -2x^2+8x+10

OpenStudy (anonymous):

the second question is x^2-2x-35

OpenStudy (anonymous):

first = -2(x^2-4x-5) = -2(x-5)(x+1)

OpenStudy (anonymous):

second = (x-7)(x+5)

OpenStudy (anonymous):

thnanx